OpenStudy (anonymous):

what is the molecular mass of a gas if 2.1 g of the gas occupies the volume of 1.72 L at 25 degrees celcius and 750 mm Hg???"

OpenStudy (aaronq):

use the ideal gas law and rearrange it to solve for molar mass, M PV=nRT n=m/M PV=mRT/M M=mRT/PV